In a quadratic equation, once you have something that looks like this-
-9(plus/minus)(square root) 121. Do you (plus minus) the square root of the number or the actual number?

Answer by stanbon(75887)

In a quadratic equation, once you have something that looks like this-
-9(plus/minus)(square root) 121. Do you (plus minus) the square root of the number or the actual number?
----
you add to get one value; then you subtract to get a second value
-----
-9 +/-sqrt(121)
---
-9+sqrt(121) = -9+11 = 2
----
-9-sqrt(121) = -9-11 = -20
==================================
Cheers,
Stan H.